OpenStudy (waheguru):

How Do I find the vertex in the porabola by the equation y=-2(x+4)(2x-5)

OpenStudy (waheguru):

I know how to find the y - intercept which is 40 but I am getting the wrong answer for the vertex

OpenStudy (anonymous):

several different ways to locate the vertex of your given equation.... 1-you can find the x-intercepts and average them out and that will give you the x-coordinate of the vertex. to find the y-coordinate just plug that x value back into the original equation. 2-you can put the equation in standard form: y = ax^2 + bx + c, and the x-coordinate of your vertex is -b/(2a). find the y-coordinate the same way as in #1. 3-complete the square and get your equation in vertex form: \(\large y-k=a(x-h)^2 \) but this will require completing the square.

OpenStudy (waheguru):

Sir we did a bit of a lesson on the first method but i did nto understand it

OpenStudy (waheguru):

How Do I find the x- co ordinate of the vertex

OpenStudy (waheguru):

One will for sure be -4 but what about the second part that has 2x-5?

OpenStudy (anonymous):

Are you punjabi

OpenStudy (anonymous):

yes... just set that equal to zero and solve for x: \(\large 2x-5=0 \)

OpenStudy (anonymous):

I'm just wondering because I'm punjabi too

OpenStudy (waheguru):

yea, thats what out teacher said to make the factors equal 0 so does that mean that the x co-ordinate of the second value is -5?

OpenStudy (waheguru):

I mean 5

OpenStudy (anonymous):

it shoulg be 5/2...

OpenStudy (anonymous):

*should

OpenStudy (anonymous):

Your write it should be 5/2

OpenStudy (waheguru):

So it would be 2.5

OpenStudy (anonymous):

Yah

OpenStudy (anonymous):

You didn't answer me are you a punjabi

OpenStudy (anonymous):

the average of the zero is -3/4.... that's what i got. so the x-coordinate of the vertex is -3/4

OpenStudy (anonymous):

can u verify that it is -3/4 ??? i'm a little bit out of it right now so my calculations may be off...

OpenStudy (waheguru):

so 4+2/5 divided by 2right?